How many liters of a 60% acid solution must be mixed with a 75% acid solution to get 20L of a 72% solution?
Answer by nerdybill(7384)

How many liters of a 60% acid solution must be mixed with a 75% acid solution to get 20L of a 72% solution?
.
Let x = liters of 60% acid solution to be added
then
20-x = liters of 75% acid solution to be added
.
Our equation:
.60x + .75(20-x) = .72(20)
.60x + 15 - .75x = 14.4
15 - .15x = 14.4
- .15x = -.6
x = 4 L (60% solution)
.
75% solution:
20-x = 20-4 = 16 L